OpenStudy (anonymous):

Find the solution to the equation -3 - 9 = -12 A. x = 7 B. x = -7 C. x = 1 D. x = -1 i keep coming up with 14...

OpenStudy (anonymous):

where si the x

OpenStudy (anonymous):

i think he forget to include x to the first term of the equation

OpenStudy (anonymous):

oops it is -3x - 9 = -12

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

got it now! i realized what i did wrong as soon as you said that :)

OpenStudy (uri):

Oh,in that case -3x=-12+9 -3x=-3 x=-3/-3 x=1

OpenStudy (anonymous):

thanks so much both of you :)

OpenStudy (anonymous):

then where is our medal... yu didnt give me medal for the first questiion i answered :/

OpenStudy (anonymous):

thanks for the medal :)

OpenStudy (anonymous):

@princed @uri thanks you two i got 97% on my practice yayyy

OpenStudy (anonymous):

its our pleasure to assist you... feel free to bring another one anytime:)